# Transformer representation learning is necessary for dynamic multi-modal physiological data on small-cohort patients

### v1

@@ 2.1 Patient characteristics and feature selection
All data were collected from patients admitted to the ICU of HOSPITAL following cardiovascular surgery. Physiological data recording commenced postoperatively once the patients were awake and continued for approximately two hours. Clinical diagnoses of POD from postoperative days 1 to 3 were determined by clinicians based on standardized delirium scale scores. Initially, data from 61 patients were collected. After excluding individuals with missing data or incomplete diagnostic records, the final study cohort comprised 56 patients. The surgical procedures included heart valve surgery (18 cases), hemi-aortic arch replacement (17 cases), coronary artery bypass grafting (9 cases), Bentall/Wheat procedures (7 cases), and total aortic arch replacement (5 cases). The average surgical duration was 326.16 minutes. Patients were further classified into two groups (TYPE I and TYPE II) based on the quality of their physiological data. TYPE I included patients whose data mostly remained within acceptable physiological ranges, whereas TYPE II comprised the remaining patients.
@@ 2.2 Performance and visualization of representation
Transformer25 $\surd$ $x_{i,:}\in\mathbb{R}^{d_{m}}$
Autoformer26 $\surd$ $x_{i,:}\in\mathbb{R}^{d_{m}}$
Informer27 $\surd$ $x_{i,:}\in\mathbb{R}^{d_{m}}$
iTransformer28 $\surd$ $\surd$ $x_{T,:}\in\mathbb{R}^{T\times d_{m}}$
Crossformer29 $\surd$ $\surd$ $x_{i:i+t,j:j+d}\in\mathbb{R}^{t\times d}$
Origin Pathformer30 $\surd$ $\surd$ $x_{i:i+t,j}\in\mathbb{R}^{t}$
@@
Table 2: Existing Transformer models utilize different parts of one time-series $X_{in}\in\mathbb{R}^{T\times d_{m}}$ as tokens.
@@
Figure 1: Visualization of training set using t-SNE dimension reduction. Positive samples came from POD patients while negative samples came from non-POD patients. We generated the samples by window sliding, the total sample size is 1824 with prediction period of 30 minutes.
@@ 2.3 Ablation and hyperparameter settings
Table 4: Ablation study conducted on fusion adaptation and TrendLoss, we only use Pathformer as an example with prediction period $T=30\text{minutes}$ and TrendLoss regularization strength $\lambda=5^{-4}$. The best results are highlighted in bold across different models, and the second-best results are underlined.
@@
Figure 2: Impact of varying TrendLoss regularization strength $\lambda$ on POD1-3 emphasizes different optimal choices for different POD indicators.
@@
Figure 3: Impact of varying prediction period $T$ emphasizes that macrospoic information and longer monitoring duration are essential for POD diagnosis.
@@ 2.4 Distinct outcomes on patient TYPE II
Given the distinct preprocessing methods applied to different patient subtypes, we assessed the efficacy of various Transformer representations under standardized settings. Transformer architectures evaluated for patient TYPE II included Transformer25, Autoformer26, Informer27, Crossformer29, iTransformer28, and Pathformer30, with detailed outcomes presented in Table 6 of the appendix. Notably, Transformer and Autoformer demonstrated significant performance improvements, particularly evident in AUROC and AUPRC across all POD indicators. For instance, the Transformer model achieved an AUROC of 0.8622 for POD3, a notable increase compared to the baseline SVM classifier. Informer similarly showed robust performance for patient TYPE II. Notably, these three Transformer models process each dimension independently, applying self-attention exclusively in the temporal dimension. In contrast, models employing patch mechanisms demonstrated increased specificity but markedly decreased sensitivity, with Fusion Pathformer yielding the lowest sensitivity. These findings suggest temporal dependencies may play a more crucial role than inter-dimensional relationships in distinguishing POD from non-POD patients in TYPE II.
@@ 3 Discussion
Observation 1: Transformer representations enhance sensitivity for POD1-3 indicators but can adversely affect specificity, particularly for POD1.
@@
As shown in Table 1, the Fusion Pathformer and iTransformer consistently improved sensitivity, while Crossformer notably improved specificity. This suggests that Crossformer’s attention mechanism, which integrates relationships across multiple dimensions, may preferentially distinguish non-POD patients by emphasizing both inter-dimensional and temporal relationships.
Observation 2: Among patch-based Transformer models, the Fusion Pathformer notably demonstrated the greatest sensitivity enhancement.
@@
Fusion/Origin PathFormer, Crossformer are three Transformers that implement patch mechanism. In the Table 1, our Fusion Pathformer significantly improved sensitivity, likely due to its enhanced capacity to effectively capture the pronounced abnormal physiological fluctuations observed in POD patients with its adaptive multi-scale router. In constrast, the fixed length segmentation in Crossformer may have obscured the critical temporal features.
Observation 3: Temporal features appear particularly crucial in identifying POD patients.
@@
Informer, which focuses exclusively on the attention over the whole time-series, often surpassed Crossformer, even sometimes fusion Pathformer in sensitivity. This may imply that temporal characteristics are essential than relationship between different dimensions in POD patients. This appears intuitive, as POD patients often exhibit noticeable changes in their physiological signals like the drastic fluctuation of heart rate(HR) as well as a slowdown in electroencephalographic activity. This suggests that in clinical monitoring, continuously tracking variations in a single-modality even one physiological signal may be more beneficial.
Observation 4: Transformer representations show preference for Logistic Regression compared to Support Vector Machine.
@@
According to Table 1, improvements in the Youden index were substantially greater for LR than for SVM following Transformer representations. These findings indicate that Transformer-derived features align better with linear classifiers such as Logistic Regression for POD prediction tasks.
Observation 5: TrendLoss enhances downstream classification performance, particularly when pre-training involves high prediction error.
@@
Comparing Fusion Pathformer trained with and without TrendLoss, the former consistently exhibited superior sensitivity, specificity and AUROC metrics when coupled with LR. Given the complexity of high-dimensional, long-term physiological data, reliance solely on mean squared error(MSE) appears insufficient, suggesting the necessity of additional regularization methods such as TrendLoss.
Observation 6: Macroscopic and coarse-grain temporal information plays crucial role in POD diagnosis.
@@
Experiments utilizing varying prediction period revealed substantial performance differences. Specifically, classification results significantly deteriorated when the prediction period was shorter than 30 minutes, emphasizing that macroscopic information is pivotal for POD diagnosis as they usually provide more information about the fluctuation and changes in the physiological state of one patients. Consequently, monitoring periods of at least 30 minutes are recommended for reliable clinical delirium assessments, aligned with another research using Transformer to predict POD on intraoperative temporal dynamics33.
Observation 7: Logistic Regression and SVM alone fail in effectively distinguish between POD and non-POD patients for TYPE II.
@@
Experiments conducted on patient TYPE II reflect that standalone LR and SVM classifiers are incapable of differentiating POD patients from non-POD patients. The extremely low sensitivity suggests that distinctions between POD and non-POD patients are not well manifested in the raw physiological signals. This highlights the necessity of advanced representation methods to capture informative patterns in such challenging patient cohorts.
Observation 8: Transformer architectures focused exclusively on temporal dependency outperform other Transformer variants for patient TYPE II.
@@
Our results show that Transformer models like Transformer, Autoformer and Informer, which apply self-attention solely in the temporal dimension and involve no information about different modalities, substantially improved both sensitivity and specificity, particularly evident for POD3. These findings suggest that for patient TYPE II, temporal fluctuations are more clinically informative than inter-dimensional relationships, although specificity enhancements observed in other Transformer variants are also noteworthy.
@@ 4.3 Framework building and module design
Figure 4: Our framework of representation-based classification comprises of three main parts, data preprocessing, representation learning and predictive classification learning.
@@
Meantime, we proposed a fusion adaptation of Pathformer30 as a new backbone. Our adaptation can be summarized as: different modalities share the same transformer-like block, namely AMS block proposed in original Pathformer to share the knowledge between each other. Router component will capture the dynamic characteristic of each modality and allocate proper Transformer blocks for each of them. This means different modalities may be assigned to different AMS blocks according to their dynamic features and utilize Patching mechanism43, 44 to partition time-series into parts.
@@ 4.3.1 Multi-modal adaptation of Pathformer
Figure 5: Fusion adaptation of Pathformer, each modality in one input time-series will be routed to topK gates each of which is connected with one Transformer block. Before inputing each modality into Transformer blocks, alignment is done to embed different modalities into the same size. Each small square in the figure corresponds to one timestamp $X_{i}^{(m)}\in\mathbb{R}^{d_{e}}$.
@@ 4.4.1 Transformer representation learning and inference
Autoregression and self-supervision are two frequently encountered tasks and methodologies in time series analysis45, 46. Autoregression, also prestigious for time series forecasting, has been extensively studied47. In recent years, researches based on the Transformer architecture has proliferated, like Autoformer26, Informer27, Crossformer29, iTransformer28, all concentrating on time series prediction with auto-regression.
We selected autoregression task for representation learning. To accomplish this goal, we add a linear decoder to project the $\bar{X}_{h}$ to $X_{pred}$.
@@ 4.4.2 Classifier training and testing
In classifier training/testing phase, to evaluate the effect of Transformer representation, we designed two different dataset preparation methods to compare the differences under identical settings with and without Transformer representations.

### v2

@@ Abstract
Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China Nanjing University]National Laboratory of Solid State Microstructure, Department of Physics, Nanjing University, Nanjing 210093, China Nanjing Drum Tower Hospital]Department of Thoracic and Cardiovascular Surgery, The Affiliated Drum Tower Hospital of Nanjing University Medical School, Kuang Yaming Honors School, Nanjing University, Nanjing 210023, China
@@ 1 Introduction
Figure 1: Overview of study design. Post-operative delirium, a prevalent acute neuropsychiatric syndrome is crucial and challenging for early diagnosis. A cascading structure of predictive system was established with components of data preprocessing, Transformer representation learning and machine learning classifier, utilizing multi-modal physiological data including aEEG, vital signs, ECG and hemodynamics. Based on Pathformer25, we proposed fusion adaptation version for multi-modal physiological data and introduced a new regularization item TrendLoss to assist with representation learning. We investigated the average prediction performance of post-operative day 1 to 3 with several Transformer representations. AUROC, sensitivity and representation visualization exhibited a profound improvement with our Fusion Pathformer.
@@ 2.1 Patient characteristics and feature selection
All data were collected from patients admitted to the ICU of The Affiliated Drum Tower Hospital of Nanjing University Medical School following cardiovascular surgery. Physiological data recording commenced postoperatively once the patients were awake and continued for approximately two hours. Clinical diagnoses of POD from postoperative days 1 to 3 were determined by clinicians based on standardized delirium scale scores. Initially, data from 61 patients were collected. After excluding individuals with missing data or incomplete diagnostic records, the final study cohort comprised 56 patients. The surgical procedures included heart valve surgery (18 cases), hemi-aortic arch replacement (17 cases), coronary artery bypass grafting (9 cases), Bentall/Wheat procedures (7 cases), and total aortic arch replacement (5 cases). The average surgical duration was 326.16 minutes. Patients were further classified into two groups (TYPE I and TYPE II) based on the quality of their physiological data. TYPE I included patients whose data mostly remained within acceptable physiological ranges, whereas TYPE II comprised the remaining patients.
@@ 2.2 Performance and visualization of representation
Transformer26 $\surd$ $x_{i,:}\in\mathbb{R}^{d_{m}}$
Autoformer27 $\surd$ $x_{i,:}\in\mathbb{R}^{d_{m}}$
Informer28 $\surd$ $x_{i,:}\in\mathbb{R}^{d_{m}}$
iTransformer29 $\surd$ $\surd$ $x_{T,:}\in\mathbb{R}^{T\times d_{m}}$
Crossformer30 $\surd$ $\surd$ $x_{i:i+t,j:j+d}\in\mathbb{R}^{t\times d}$
Origin Pathformer25 $\surd$ $\surd$ $x_{i:i+t,j}\in\mathbb{R}^{t}$
@@
Table 2: Token unit comparison between several competitive Transformer architectures. Existing Transformer models utilize different parts of one time-series $X_{in}\in\mathbb{R}^{T\times d_{m}}$ as tokens.
@@
Figure 2: Visualization of training set using t-SNE dimension reduction. Positive samples came from patients diagnosed POD on the first post-operative day while negative samples came from those diagnosed normal. We generated the samples by window sliding, the total sample size is 1824 with prediction period of 30 minutes.
@@ 2.3 Ablation and hyperparameter settings
Table 4: Ablation study conducted on fusion adaptation and TrendLoss. We only use Pathformer as an example with prediction period $T=30\text{ minutes}$ and TrendLoss regularization strength $\lambda=5^{-4}$. The best results are highlighted in bold across different models, and the second-best results are underlined.
@@
Figure 3: Impact of TrendLoss regularization strength $\lambda$. Different results on POD1-3 emphasizes different optimal choices for different POD indicators and different machine learning classifiers.
@@
Figure 4: Impact of varying prediction period $T$. Results of Logistic Regression and SVM under all POD indicators exhibit the similar pattern when prediction period $T$ is changing. Subfigure (A)(B)(C) indicate that the diagnosis of POD patients is quite sensitive to the prediction period $T$. These results emphasize that macrospoic information and longer monitoring duration are critical for POD diagnosis.
@@ 2.4 Distinct outcomes on patient TYPE II
Given the distinct preprocessing methods applied to different patient subtypes, we assessed the efficacy of various Transformer representations under standardized settings. Transformer architectures evaluated for patient TYPE II included Transformer26, Autoformer27, Informer28, Crossformer30, iTransformer29, and Pathformer25, with detailed outcomes presented in Table 6 of the appendix. Notably, Transformer and Autoformer demonstrated significant performance improvements, particularly evident in AUROC and AUPRC across all POD indicators. For instance, the Transformer model achieved an AUROC of 0.8622 for POD3, a notable increase compared to the baseline SVM classifier. Informer similarly showed robust performance for patient TYPE II. Notably, these three Transformer models process each dimension independently, applying self-attention exclusively in the temporal dimension. In contrast, models employing patch mechanisms demonstrated increased specificity but markedly decreased sensitivity, with Fusion Pathformer yielding the lowest sensitivity. These findings suggest temporal dependencies may play a more crucial role than inter-dimensional relationships in distinguishing POD from non-POD patients in TYPE II.
@@ 3 Discussion
Transformer representations enhance sensitivity for POD1-3 indicators but can adversely affect specificity, particularly for POD1. As shown in Table 1, the Fusion Pathformer and iTransformer consistently improved sensitivity, while Crossformer notably improved specificity. This suggests that Crossformer’s attention mechanism, which integrates relationships across multiple dimensions, may preferentially distinguish non-POD patients by emphasizing both inter-dimensional and temporal relationships.
@@
Among patch-based Transformer models, the Fusion Pathformer notably demonstrated the greatest sensitivity enhancement. Fusion/Origin PathFormer, Crossformer are three Transformers that implement patch mechanism. In the Table 1, our Fusion Pathformer significantly improved sensitivity, likely due to its enhanced capacity to effectively capture the pronounced abnormal physiological fluctuations observed in POD patients with its adaptive multi-scale router. In constrast, the fixed length segmentation in Crossformer may have obscured the critical temporal features.
@@
Temporal features appear particularly crucial in identifying POD patients. Informer, which focuses exclusively on the attention over the whole time-series, often surpassed Crossformer, even sometimes fusion Pathformer in sensitivity. This may imply that temporal characteristics are essential than relationship between different dimensions in POD patients. This appears intuitive, as POD patients often exhibit noticeable changes in their physiological signals like the drastic fluctuation of heart rate(HR) as well as a slowdown in electroencephalographic activity. This suggests that in clinical monitoring, continuously tracking variations in a single-modality even one physiological signal may be more beneficial.
@@
Transformer representations show preference for Logistic Regression compared to Support Vector Machine. According to Table 1, improvements in the Youden index were substantially greater for LR than for SVM following Transformer representations. These findings indicate that Transformer-derived features align better with linear classifiers such as Logistic Regression for POD prediction tasks.
@@
TrendLoss enhances downstream classification performance, particularly when pre-training involves high prediction error. Comparing Fusion Pathformer trained with and without TrendLoss, the former consistently exhibited superior sensitivity, specificity and AUROC metrics when coupled with LR. Given the complexity of high-dimensional, long-term physiological data, reliance solely on mean squared error(MSE) appears insufficient, suggesting the necessity of additional regularization methods such as TrendLoss.
@@
Macroscopic and coarse-grain temporal information plays crucial role in POD diagnosis. Experiments utilizing varying prediction period revealed substantial performance differences. Specifically, classification results significantly deteriorated when the prediction period was shorter than 30 minutes, emphasizing that macroscopic information is pivotal for POD diagnosis as they usually provide more information about the fluctuation and changes in the physiological state of one patients. Consequently, monitoring periods of at least 30 minutes are recommended for reliable clinical delirium assessments, aligned with another research using Transformer to predict POD on intraoperative temporal dynamics33.
@@
Logistic Regression and SVM alone fail in effectively distinguish between POD and non-POD patients for TYPE II. Experiments conducted on patient TYPE II reflect that standalone LR and SVM classifiers are incapable of differentiating POD patients from non-POD patients. The extremely low sensitivity suggests that distinctions between POD and non-POD patients are not well manifested in the raw physiological signals. This highlights the necessity of advanced representation methods to capture informative patterns in such challenging patient cohorts.
@@
Transformer architectures focused exclusively on temporal dependency outperform other Transformer variants for patient TYPE II. Our results show that Transformer models like Transformer, Autoformer and Informer, which apply self-attention solely in the temporal dimension and involve no information about different modalities, substantially improved both sensitivity and specificity, particularly evident for POD3. These findings suggest that for patient TYPE II, temporal fluctuations are more clinically informative than inter-dimensional relationships, although specificity enhancements observed in other Transformer variants are also noteworthy.
@@ 4.3 Framework building and module design
Figure 5: Our design of representation-based post-operative delirium prediction model. The model comprises three main parts, data preprocessing, representation learning and predictive classification learning.
@@
Meantime, we proposed a fusion adaptation of Pathformer25 as a new backbone. Our adaptation can be summarized as: different modalities share the same transformer-like block, namely AMS block proposed in original Pathformer to share the knowledge between each other. Router component will capture the dynamic characteristic of each modality and allocate proper Transformer blocks for each of them. This means different modalities may be assigned to different AMS blocks according to their dynamic features and utilize Patching mechanism43, 44 to partition time-series into parts.
@@ 4.3.1 Multi-modal adaptation of Pathformer
Figure 6: Fusion adaptation of Pathformer. Each modality in one input time-series will be routed to topK gates each of which is connected with one Transformer block. Before inputting each modality into Transformer blocks, alignment is done to embed different modalities into the same size. Each small square in the figure corresponds to one timestamp $X_{i}^{(m)}\in\mathbb{R}^{d_{e}}$.
@@ 4.4.1 Transformer representation learning and inference
Auto-regression and self-supervision are two frequently encountered tasks and methodologies in time series analysis45, 46. Auto-regression, also prestigious for time series forecasting, has been extensively studied47. In recent years, researches based on the Transformer architecture has proliferated, like Autoformer27, Informer28, Crossformer30, iTransformer29, all concentrating on time series prediction with auto-regression.
We selected auto-regression task for representation learning. To accomplish this goal, we add a linear decoder to project the $\bar{X}_{h}$ to $X_{pred}$.
@@ 4.4.2 Classifier training and testing
In classifier training and testing phase, to evaluate the effect of Transformer representation, we designed two different dataset preparation methods to compare the differences under identical settings with and without Transformer representations.